OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1). (4 points)parallel: y = -x - 1 perpendicular: y = x + 2 parallel: y = x - 1 perpendicular: y = -x + 1 parallel: y = x + 1 perpendicular: y = -x - 3 parallel: y = 2x - 2 perpendicular: y = -2x - 1

OpenStudy (anonymous):

You can graph each equation by filling in values then look and see if the lines drawn are parallel or perpendicular or if they pass through point (-2,-1).

OpenStudy (anonymous):

parallel: y = x + 1 perpendicular: y = -x - 3